OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20 x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution

Directrix (directrix):

|2x + 6| − 4 = 20 Begin by adding 4 to both sides of the equation. Do not mess with the absolute value quantity just yet.

OpenStudy (anonymous):

ok so now it [2x+6]=24

Directrix (directrix):

Good.

OpenStudy (anonymous):

now what do i do

Directrix (directrix):

By definition of an absolute value equality, [2x+6] = 24 means 2x + 6 = 24 OR 2x + 6 = -24 So, solve each of those linear equations for x.

OpenStudy (anonymous):

oh ok thanks lol :)

Directrix (directrix):

Go ahead and post the answers you get and we'll check each other.

Directrix (directrix):

@Madisons_xo_beauty

OpenStudy (anonymous):

ok one minute

OpenStudy (anonymous):

i got c

Directrix (directrix):

Take your time. Fast is fine but accuracy is everything. ~Wyatt Earp

Directrix (directrix):

I got 9 and - 15. Is that c?

Directrix (directrix):

I think it is.

OpenStudy (anonymous):

lol yes

OpenStudy (anonymous):

thankyou! @Directrix

Directrix (directrix):

You are welcome. Thanks for being an active participant in this problem.

OpenStudy (anonymous):

no problem